OpenStudy (perl):

Algebra Question: $500 is invested in a stock that returns 12%. How much of an additional $3200 should be invested in each of two different stocks earning 18% and 10% in order to earn a total return of 16%.

OpenStudy (perl):

i can't seem to get the books answer ,

OpenStudy (perl):

book says $1,650 at 18% and $1,550 at 10%

OpenStudy (perl):

I did .18( x) + .10 (y ) + .12 * 500 = .16 ( 3200 + 500) x + y = 3200

OpenStudy (kropot72):

The book's answer is incorrect. Correct answers are $2650 at 18% and $550 at 10%.

OpenStudy (perl):

something is very odd here

OpenStudy (perl):

help anyone

OpenStudy (perl):

hi krotom :)

OpenStudy (perl):

I plugged the two equations into wolfram, and i got the books answer . or do you disagree with my equation set up

OpenStudy (kropot72):

Your equation setup in Wolf is different from your posted set up. In wolf you have the total return as 0.16 * 3200. However your posting has " .16 ( 3200 + 500)"

OpenStudy (perl):

right, i assumed we should add 500 , because it says 'total return'

OpenStudy (perl):

$500 is invested in a stock that returns 12%. How much of an additional $3200 should be invested in each of two different stocks earning 18% and 10% in ***order to earn a total return of 16%.***

OpenStudy (kropot72):

I think that the correct interpretation of the question is to take the total investment as being $500 + $3200. If otherwise, why mention the $500 and its return?

OpenStudy (perl):

ok then you agree with the book's answer?

OpenStudy (perl):

right, it seems odd to mention it otherwise (if we don't include it)

OpenStudy (kropot72):

No, I disagree with the book's answer.

OpenStudy (perl):

book says $1,650 at 18% and $1,550 at 10%

OpenStudy (perl):

what was your equation set up?

OpenStudy (kropot72):

My equations are basically the same as yours posted here. The first equation on Wolf states the return on three investments on the left hand side, but has the return on only two investments on the right hand side. Therefore that equation on Wolf is invalid.

OpenStudy (perl):

ok so you disagree with this .18( x) + .10 (y ) + .12 * 500 = .16 ( 3200 + 500)

OpenStudy (kropot72):

".18( x) + .10 (y ) + .12 * 500 = .16 ( 3200 + 500) " I agree with the above equation. The above is not what you have on Wolf. This is what is on Wolf (as below): .18( x) + .10 (y ) + .12 * 500 = .16 ( 3200) I disagree with that.

OpenStudy (perl):

wow, i feel stupid i typed a new entry into wolfram, here it is fixed http://www.wolframalpha.com/input/?i=.18x%2B.10y+%2B+.12*500+%3D+.16%283200%2B500%29%2C+x+%2B+y+%3D+3200

OpenStudy (kropot72):

np. Sorry I wasn't clear enough earlier. :)

OpenStudy (perl):

darn, this is strange when the book has such an error

OpenStudy (perl):

i tried to reverse engineer, the only way to get the books answer was to remove the 500 part

OpenStudy (perl):

this equation will give you the book's answer (but the equation seems wrong) http://www.wolframalpha.com/input/?i=.18x%2B.10y+%2B+.12*500+%3D+.16%283200%29%2C+x+%2B+y+%3D+3200

OpenStudy (perl):

i really hate book mistakes , thats one of my pet peeves

OpenStudy (kropot72):

I think you have probably found the error made when the book's answer was calculated. Answers in textbooks should be thoroughly and independently checked. Purchasers are getting poor value for their money when answers in books can't be trusted.

OpenStudy (perl):

i agree